Chapter 30, Problem 066 A circular loop of wire 82 mm in radius carries a current of 115 A. Find the (a) magnetic field strength and (b) energy density at the center of the loop. (a) Number Units (b) Number Units
Answer #1

Solution :-

Given that

radius (R)= 82mm

Current( I)=115 A

a)

for magnetic field strength

B=μoI/2R

Where μo =4πx10-7 T/mA

B=(4πx10-7x115)/(2x82)

= 8.807 x10^-7 T

b)

energy density

ub=1/2B2o

      = μoI2 /8R2

=(4πx10-7x(115)2)/(8x82x82)

=3.0879x10-7 J/m3